A=$6000,r=8,t=2years
Found 2 solutions by Mathtut, checkley77:
Answer by Mathtut(3670)

A=P(1+rt)
:
6000=P(1+.08(2))
:
6000=1.16P
:
P=$5172.41

Answer by checkley77(12844)

A=$6000,r=8,t=2years
A(1+R)^T
6000(1+.08)^2
6000(1.08)^2
6000*1.1664=$6,998.40 ANS.